# Extended Galerkin neural network approximation of singular variational problems with error control

### v1

@@ 1 Introduction
where $\mathcal{L}$ and $\mathcal{B}$ are linear differential operators. The only assumption needed is the the rather mild condition that the problem is well-posed in the sense that it admits a unique solution that depends continuously on the data using appropriate norms. Under this condition, we show that our previous results extend to the more general scenario given in (1) including the provision of a computable a posteriori error estimator based on using least square formulation. While least squares formulations have featured in previous work using neural networks to approximate PDEs in a somewhat ad hoc fashion – e.g. mean-squared error (MSE) or $L^{2}$ norm as a loss metric – we argue that the least squares formulation is dictated by the properties of the underlying continuous problem (1) rather than the whim of the user if one wishes to obtain a rigorous convergence theory. Ad hoc choices of formulation correspond to a BVP which may not be uniquely solvable or for which there is no continuous dependence on the data. We shall demonstrate in this work that the use of such formulations can result in nonphysical solutions (see Examples 3.2 and 3.5)
@@ 2 Extended Galerkin neural networks (xGNN)
We first give a brief summary of the basic Galerkin neural network framework for symmetric and positive-definite problems in Section 2.1 following [1]. We then demonstrate how to rigorously extend this approach to non-symmetric and/or indefinite and negative-definite problems in Section 2.2 as well as enrich the neural network approximation space with known solution structures in Section 2.3. Collectively, we refer to these advancements as extended Galerkin neural networks (xGNN).
@@ 2.1 The basic Galerkin neural network framework
where $a$ is assumed to be a symmetric positive-definite, continuous, and coercive bilinear form with respect to $(X,||\cdot||_{X})$ which defines an energy norm $|||\cdot|||:=a(\cdot,\cdot)^{1/2}$, and $F$ is assumed to be continuous. Given an initial approximation $u_{0}\in X$ to (2), the goal of the Galerkin neural network is to iteratively construct a finite-dimensional subspace $S_{j}:=\text{span}\{u_{0},\varphi_{1}^{NN},\dots,\varphi_{j}^{NN}\}$ such that the basis function $\varphi_{i}^{NN}$ satisfies
@@
is the weak residual in the current approximation $u_{i-1}$, $B$ is the closed unit ball in $X$, and $V_{\mathbf{n},L}^{\sigma}$, $\mathbf{n}^{(i)}\in\mathbb{N}^{L}$ is the set
$V_{\mathbf{n},L}^{\sigma}:=$$\left\{v\in X\;:\;v(x)=(\sigma\circ\mathbf{T}_{L}\circ\dots\sigma\circ\mathbf{T}_{1}(x))\cdot\mathbf{c},\;\mathbf{T}_{i}(t):=t\cdot\mathbf{W}^{(i)}+\mathbf{b}^{(i)}\right.$
(4) $\left.\;\mathbf{W}^{(i)}\in\mathbb{R}^{n_{i-1}\times n_{i}},\;\mathbf{b}^{(i)}\in\mathbb{R}^{1\times n_{i}},\;\mathbf{c}\in\mathbb{R}^{n_{L}\times 1},\;x\in\mathbb{R}^{1\times d}\right\},$
with the convention that the data $x$ has dimension $\mathbb{R}^{1\times d}$ and $n_{0}=d$. The set $V_{\mathbf{n},L}^{\sigma}$ describes the realizations of a multilayer feedforward neural network with depth $L$, widths $\mathbf{n}$, and activation function $\sigma$. The set $V_{\mathbf{n}^{(i)},L_{i}}^{\sigma,C}$ denotes functions in $V_{\mathbf{n}^{(i)},L_{i}}^{\sigma}$ with bounded parameters:
(5) $V_{\mathbf{n},L}^{\sigma,C}:=\left\{v\in V_{\mathbf{n},L}^{\sigma}:\sum_{i=1}^{L}||\mathbf{W}^{(i)}||_{\infty}+||\mathbf{b}^{(i)}||_{\infty}+||\mathbf{c}||_{\infty}\leqslant C\right\}.$
@@
(7) $\begin{dcases}\mathbf{A}\mathbf{c}&=\mathbf{F}\\ \mathbf{A}_{jk}&=a(\sigma(x\cdot\mathbf{W}_{j}+\mathbf{b}_{j}),\sigma(x\cdot\mathbf{W}_{k}+\mathbf{b}_{k}))\\ \mathbf{F}_{j}&=L(\sigma(x\cdot\mathbf{W}_{j}+\mathbf{b}_{j}))-a(u_{i-1},\sigma(x\cdot\mathbf{W}_{k}+\mathbf{b}_{k}))\end{dcases}$
in order to update the activation coefficients $\mathbf{c}$. For ease of notation, we consider only the case when $L=1$ in (7) while noting that the linear system for $L>1$ differs only in that $\sigma(x\cdot\mathbf{W}_{j}+\mathbf{b}_{j})$ will be replaced by the corresponding more complicated expression for the $j$th component of $\sigma\circ\mathbf{T}_{L}\circ\dots\sigma\circ\mathbf{T}_{1}(x)$. The notation $\mathbf{W}_{j}$ and $\mathbf{b}_{j}$ are shorthand for $\mathbf{W}[:,j]$ and $\mathbf{b}[1,j]$, respectively. The linear system in (7) corresponds to the orthogonal projection of the error $u-u_{i-1}$ onto the subspace $\Phi:=\text{span}\{\sigma(x\cdot\mathbf{W}_{j}+\mathbf{b}_{j})\}_{j=1}^{n}$.
@@ 2.2 Extension to non-self-adjoint and non-positive-definite problems
The framework described in Section 2.1 considers self-adjoint, positive definite boundary value problems whose variational formulation naturally gives rise to a bilinear form $a(\cdot,\cdot)$ that is symmetric positive-definite. Here, we extend the approach to a more general boundary value problem with the strong form
@@
Variational formulations such as (13) are employed in least squares finite element methods [7] often with a reduction of the PDE to first order. In the current work, we make no such conversion to first-order systems but note that this conversion is fully-compatible with the proposed method. The operator $a_{LS}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is symmetric positive-definite regardless of the structure of $\mathcal{L}$. Moreover, if the operators $\mathcal{L}$ and $\mathcal{B}$ satisfy the above assumptions, then the boundary value problem (11) can be expressed in the variational form (13):
@@
(17) $||\mathbf{u}||_{\mathbf{X}}\leqslant C_{3}(||\mathbf{f}||_{\mathcal{V}}+||\mathbf{g}||_{\mathcal{W}})$
for some constant $C_{3}>0$. Then the bilinear form $a_{LS}$ is continuous and coercive, and the linear form $L_{LS}$ is continuous in the sense that there exist constants $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}>0$ such that
@@
$||\mathbf{v}||_{\mathcal{X}}^{2}\leqslant 2C_{3}^{2}(||\mathcal{L}[\mathbf{v}]||_{\mathcal{V}}^{2}+||\mathcal{B}[\mathbf{v}]||_{\mathcal{W}}^{2})\leqslant 2C_{3}^{2}a_{LS}(\mathbf{v},\mathbf{v}).$
Figure 1: Errors with respect to the number of basis functions for $\delta=1$, $\mathcal{W}=L^{2}(\partial\Omega)$, and the sequence of problems $m=2,4,8,16$.
Example 2.3.
Consider a Poisson equation with exact solution given by the harmonic function $u_{m}(r,\theta)=r^{m}\sin(m\theta)$ in the unit circle. For a fixed $m\in\mathbb{N}$, a straightforward computation reveals that
$||u_{m}||_{H^{2}(\Omega)}\propto m^{3/2},\;\;\;||u_{m}||_{H^{s}(\partial\Omega)}\propto m^{s},$
and thus we have the ratio
(19) $\delta||u_{m}||_{H^{s}(\partial\Omega)}/||u_{m}||_{H^{2}(\Omega)}\propto\delta m^{s-3/2}.$
The case $\mathcal{W}=L^{2}(\partial\Omega)$ of the boundary penalization in (12) corresponds to $s=0$, and $m^{-3/2}\to 0$ as $m\to\infty$ which implies that the constant $C_{2}$ in Theorem 2.1 becomes vanishingly small. Indeed, the estimate (12) for the Poisson equation reads [11]
$||u||_{H^{2}(\Omega)}\leqslant C(||f||_{L^{2}(\Omega)}+||g||_{H^{3/2}(\partial\Omega)}),$
which indicates that the $H^{2}(\Omega)$ formulation of the Poisson equation based on (13) requires the choices $\mathcal{W}=H^{3/2}(\partial\Omega)$, $\mathcal{V}=L^{2}(\Omega)$, and $\mathcal{X}=H^{2}(\Omega)$.
One practical problem with choosing $\mathcal{W}$ to be of fractional order is the difficulty of evaluating the $H^{3/2}$-norm. For the present example where the true solution has the form $r^{m}\sin(m\theta)$, we take advantage of (19) and replace the $H^{3/2}$-norm by a weighted $L^{2}$-norm, i.e. $||u_{m}||_{H^{3/2}(\partial\Omega)}=m^{3/2}||u_{m}||_{L^{2}(\partial\Omega)}$. This is equivalent to taking $\mathcal{W}=L^{2}(\partial\Omega)$ and penalty parameter $\delta=(m^{3/2})^{2}=m^{3}$. Of course, this approach will not be possible in general but will suffice for the current example.
We first fix $\delta=1$, which corresponds to the case of $L^{2}$-penalization of the boundary conditions, and consider the sequence of PDEs corresponding to $m=2,4,8,16$ in order to demonstrate the gradual loss of coercivity for this problem. Figure 1 shows the error in various metrics for the Galerkin neural network method with formulation (13). For $m=2,4$ the $L^{2}$ boundary penalization is sufficient since $||u_{m}||_{L^{2}(\partial\Omega)},||u_{m}||_{H^{3/2}(\partial\Omega)}=\mathcal{O}(1)$. However, for larger $m$ we begin to observe reduced coercivity and the rate of convergence of the errors in all metrics.
Figure 2: Errors with respect to the number of basis functions for several choices of penalty parameter $\delta=m^{\alpha}$.
Next, we fix $m=16$ and consider the choice $\delta=m^{\alpha}$ for $\alpha=0,1,2,3,4,5$ in order to demonstrate the effect of boundary penalization in different norms. Figure 2 shows the analogous results for this set of experiments. We observe that the error in the energy norm $|||\cdot|||:=a_{LS}(\cdot,\cdot)^{1/2}$ decreases monotonically regardless of the choice of $\alpha$ due to the guarantee of $u_{i}$ being the best approximation to $u$ from $S_{i}$. For $\alpha=0,1$ ($L^{2}(\partial\Omega)$ and approximate $H^{1/2}(\partial\Omega)$ boundary penalization, respectively), the errors in every metric are slow to converge. It is in this regime that the loss of coercivity is strongest. Rates improve for $\alpha=2$ (approximate $H^{1}(\partial\Omega)$ boundary penalization) and are optimal for $\alpha=3$ (approximate $H^{3/2}(\partial\Omega)$ boundary penalization). On the other hand, the overpenalization corresponding to $\alpha=4,5$ (approximate $H^{2}(\partial\Omega)$ and $H^{5/2}(\partial\Omega)$ boundary penalization, respectively) results in a loss of convergence and in particular, the clear loss of norm equivalence between ||||||||$\cdot$|||| | | ⋅ | | | and $||\cdot||_{H^{k}(\Omega)},k\in\mathbb{N}$.
The above example illustrates the importance of using the correct norms that arise in the well-posedness of the problem. In practice, the presence of fractional norms presents difficulties but is possible using techniques such as those in [19]. However, generally speaking their computation is impractical and we instead utilize the highest possible integer Sobolev norm whose order is close to the order of the desired Sobolev norm. For instance, the $H^{1}(\partial\Omega)$ penalization in Figure 2 still performs quite well when compared to the optimal $H^{3/2}(\partial\Omega)$ penalization.
2.3 Extended neural network architectures
In what follows, we assume that the solution $u\in\mathcal{X}$ admits a representation of the form
(20) $u(x)=u_{\infty}(x)+u_{\Psi}(x;\lambda),\;\;\;\;\;u_{\Psi}(x;\lambda):=\sum_{i=1}^{m}c_{\lambda,i}\Psi(x;\lambda_{i}),$
where $u_{\infty}$ is the portion of the solution with high regularity and $u_{\Psi}$ the portion with low regularity of the form $\sum_{i}c_{\Psi,i}\Psi(\cdot;\lambda_{i})$. The solutions of many classical PDEs exhibit well-known structures $\Psi(x;\lambda_{i})$, where $\lambda$ is a parameter that depends on the specific form of the domain, boundary conditions, and coefficients of the PDE. For instance, for the Poisson equation, the singular part of the solution near corners of a polygonal domain is a sum of terms of the form $\Psi(r,\theta;\lambda)\propto r^{\lambda_{i}}\sin(\lambda_{i}\theta)$, where $\lambda_{i}$ depends on the interior angle at the corner. If such knowledge-based functions are available, then it makes sense for them to be incorporated into the Galerkin neural network framework. We shall exhibit the benefits of doing so in Section 3. The extended Galerkin neural network approach we shall describe in this section gives the flexibility to incorporate this problem-specific knowledge into the basic Galerkin neural neural method.
Figure 3: Extended neural network architecture for extended Galerkin neural networks.
For a given $\Phi\in\mathcal{X}$, let
(21) $V_{m}^{\Phi}:=\{v\;:\;v(x)=\sum_{i=1}^{m}c_{\lambda,i}\Phi(x;\lambda_{i}),\;\mathbf{c}_{\lambda}\in\mathbb{R}^{m}\},$
and, as in the basic Galerkin neural network in Section 2.1, we define the set of realizations whose parameters are bounded:
(22) $V_{m}^{\Phi,C}:=\{v\in\tilde{V}_{m}^{\Phi}\;:\;||(\lambda,\mathbf{c}_{\lambda})||_{\Lambda}:=||\mathbf{c}_{\lambda}||_{\infty}+||\lambda||_{\infty}\leqslant C\}.$
We then apply the Galerkin neural network framework in Section 2.1 using the augmented space $V_{\mathbf{n},L}^{\sigma,C}\oplus V_{n_{\lambda}}^{\Phi,C}$ and seek
@@ 2.3 Extended neural network architectures
(23) $\varphi_{i}^{NN}=\operatorname*{arg\,max}_{v\in V_{\mathbf{n},L}^{\sigma,C_{1}}\bigoplus V_{n_{\lambda}}^{\Phi,C_{2}}\cap B}\langle r(u_{i-1}),v\rangle.$
In order to ensure the existence of $\varphi_{i}^{NN}$, we require that the choice of $\Phi$ results in a continuous mapping from the parameter space $(\lambda,\mathbf{c}_{\lambda})$ into the set $V_{m}^{\Phi,C}$.

### v2

@@ 1 Introduction
where $\mathcal{L}$ and $\mathcal{B}$ are linear differential operators. The only assumption needed is the the rather mild condition that the problem is well-posed in the sense that it admits a unique solution that depends continuously on the data using appropriate norms. Under this condition, we show that our previous results extend to the more general scenario given in (1) including the provision of a computable a posteriori error estimator based on using a least squares formulation. While least squares formulations have featured in previous work using neural networks to approximate PDEs in a somewhat ad hoc fashion – e.g. mean-squared error (MSE) or $L^{2}$ norm as a loss metric – we argue that the least squares formulation is dictated by the properties of the underlying continuous problem (1) rather than the whim of the user if one wishes to obtain a rigorous convergence theory. Ad hoc choices of formulation correspond to a BVP which may not be uniquely solvable or for which there is no continuous dependence on the data. We shall demonstrate in this work that the use of such formulations can result in nonphysical solutions (see Examples 3.2 and 3.5)
@@ 2 Extended Galerkin neural networks (xGNN)
We first give a brief summary of the basic Galerkin neural network framework for symmetric, positive-definite (SPD) problems in Section 2.1 following [1]. We then demonstrate how to rigorously extend this approach to non-symmetric and/or indefinite and negative-definite problems in Section 2.2 as well as enrich the neural network approximation space with known solution structures in Section 2.3. Collectively, we refer to these advancements as extended Galerkin neural networks (xGNN).
@@ 2.1 The basic Galerkin neural network framework
where $a$ is assumed to be an SPD, continuous, and coercive bilinear form with respect to $(X,||\cdot||_{X})$ which defines an energy norm $|||\cdot|||:=a(\cdot,\cdot)^{1/2}$, and $F$ is assumed to be continuous. Given an initial approximation $u_{0}\in X$ to (2), the goal of the Galerkin neural network is to iteratively construct a finite-dimensional subspace $S_{j}:=\text{span}\{u_{0},\varphi_{1}^{NN},\dots,\varphi_{j}^{NN}\}$ such that the basis function $\varphi_{i}^{NN}$ satisfies
@@
is the weak residual in the current approximation $u_{i-1}$, $B$ is the closed unit ball in $X$, and $V_{\mathbf{n},L}^{\sigma}$, $\mathbf{n}\in\mathbb{N}^{L}$ is the set
$V_{\mathbf{n},L}^{\sigma}:=$$\left\{v\in X\;:\;v(x)=(\sigma\circ\mathbf{T}_{L}\circ\dots\sigma\circ\mathbf{T}_{1}(x))\cdot\mathbf{c},\;\mathbf{T}_{j}(t):=t\cdot\mathbf{W}^{(j)}+\mathbf{b}^{(j)}\right.$
(4) $\left.\;\mathbf{W}^{(j)}\in\mathbb{R}^{n_{j-1}\times n_{j}},\;\mathbf{b}^{(j)}\in\mathbb{R}^{1\times n_{j}},\;\mathbf{c}\in\mathbb{R}^{n_{L}\times 1},\;x\in\mathbb{R}^{1\times d}\right\},$
with the convention that the data $x$ has dimension $\mathbb{R}^{1\times d}$ and $n_{0}=d$. The set $V_{\mathbf{n},L}^{\sigma}$ describes the realizations of a multilayer feedforward neural network with depth $L$, widths $\mathbf{n}$, and suitable (non-polynomial) activation function $\sigma\in X$. The set $V_{\mathbf{n},L}^{\sigma,C}$ denotes functions in $V_{\mathbf{n},L}^{\sigma}$ with bounded parameters:
(5) $V_{\mathbf{n},L}^{\sigma,C}:=\left\{v\in V_{\mathbf{n},L}^{\sigma}:||(\mathbf{W},\mathbf{b},\mathbf{c})||_{\mathcal{NN}}:=\sum_{j=1}^{L}||\mathbf{W}^{(j)}||_{\infty}+||\mathbf{b}^{(j)}||_{\infty}+||\mathbf{c}||_{\infty}\leqslant C\right\}.$
@@
(7) $\begin{dcases}\mathbf{A}\mathbf{c}&=\mathbf{F}\\ \mathbf{A}_{k\ell}&=a(\sigma(x\cdot\mathbf{W}_{k}+\mathbf{b}_{k}),\sigma(x\cdot\mathbf{W}_{\ell}+\mathbf{b}_{\ell}))\\ \mathbf{F}_{k}&=L(\sigma(x\cdot\mathbf{W}_{k}+\mathbf{b}_{k}))-a(u_{i-1},\sigma(x\cdot\mathbf{W}_{k}+\mathbf{b}_{k}))\end{dcases}$
in order to update the activation coefficients $\mathbf{c}$. For ease of notation, we consider only the case when $L=1$ in (7) while noting that the linear system for $L>1$ differs only in that $\sigma(x\cdot\mathbf{W}_{k}+\mathbf{b}_{k})$ will be replaced by the corresponding more complicated expression for the $j$th component of $\sigma\circ\mathbf{T}_{L}\circ\dots\sigma\circ\mathbf{T}_{1}(x)$. The notation $\mathbf{W}_{k}$ and $\mathbf{b}_{k}$ are shorthand for $\mathbf{W}[:,j]$ and $\mathbf{b}[1,j]$, respectively. The linear system in (7) corresponds to the orthogonal projection of the error $u-u_{i-1}$ onto the subspace $\Phi:=\text{span}\{\sigma(x\cdot\mathbf{W}_{k}+\mathbf{b}_{k})\}_{k=1}^{n}$.
@@ 2.2 Extension to non-self-adjoint and non-positive-definite problems
The framework described in Section 2.1 considers self-adjoint, positive-definite boundary value problems whose variational formulation naturally gives rise to a bilinear form $a(\cdot,\cdot)$ that is SPD. Here, we extend the approach to a more general boundary value problem with the strong form
@@
Variational formulations such as (13) are employed in least squares finite element methods [7] often with a reduction of the PDE to first order. In the current work, we make no such conversion to first-order systems but note that this conversion is fully-compatible with the proposed method.
One advantage of least squares formulations is that the bilinear form $a_{LS}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is SPD. This follows since $a_{LS}(\mathbf{v},\mathbf{v})\geqslant 0$ with equality if and only if $\mathcal{L}[\mathbf{v}]=\mathbf{0}$ and $\mathcal{B}[\mathbf{v}]=\mathbf{0}$ or, by the assumption of the uniqueness of the solution of (11), $\mathbf{v}=\mathbf{0}.$ Moreover, the variational problem (13) is well-posed:
@@
(17) $||\mathbf{u}||_{\mathcal{X}}\leqslant C_{3}(||\mathbf{f}||_{\mathcal{V}}+||\mathbf{g}||_{\mathcal{W}})$
for some constant $C_{3}>0$. Then the bilinear form $a_{LS}$ is continuous and coercive, and the linear form $F_{LS}$ is continuous in the sense that there exist constants $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}>0$ such that
@@
Similarly, from (17) we obtain $||\mathbf{v}||_{\mathcal{X}}\leqslant C(||\mathcal{L}[\mathbf{v}]||_{\mathcal{V}}+||\mathcal{B}[\mathbf{v}]||_{\mathcal{W}})$ for $\mathbf{v}\in\mathcal{X}$ and, thanks to Young’s inequality for products, deduce that the bilinear form is coercive:
$||\mathbf{v}||_{\mathcal{X}}^{2}\leqslant C_{3}^{2}(||\mathcal{L}[\mathbf{v}]||_{\mathcal{V}}+||\mathcal{B}[\mathbf{v}]||_{\mathcal{W}})^{2}\leqslant 2C_{3}^{2}(||\mathcal{L}[\mathbf{v}]||_{\mathcal{V}}^{2}+||\mathcal{B}[\mathbf{v}]||_{\mathcal{W}}^{2})\leqslant 2C_{3}^{2}a_{LS}(\mathbf{v},\mathbf{v}).$
Figure 1: Errors with respect to the number of basis functions learning using the Galerkin neural network approach applied to the variational problem (19) for $\delta=1$, $\mathcal{W}=L^{2}(\partial\Omega)$, and the sequence of problems $m=2,4,8,16$.
Example 2.3.
Consider a Poisson equation with Dirichlet data so that the corresponding least squares variational problem is given by
(19) $u\in\mathcal{X}\;:\;a_{LS}(u,v)=(-\Delta u,-\Delta v)_{\mathcal{V}}+\delta(u,v)_{\mathcal{W}}=F_{LS}(v)\;\;\;\forall v\in\mathcal{X},$
where the linear form $F_{LS}$ is chosen so that the true solution is given by the harmonic function $u_{m}(r,\theta)=r^{m}\sin(m\theta)$ for fixed $m\in\mathbb{N}$ in the unit circle.
Here, $\mathcal{X}=H^{2}(\Omega)$, $\mathcal{V}=L^{2}(\Omega)$, and $\mathcal{W}$ should be taken as $H^{3/2}(\partial\Omega)$ as suggested by the following estimate for the Poisson equation which reads [12]
(20) $||u||_{H^{2}(\Omega)}\leqslant C(||f||_{L^{2}(\Omega)}+||g||_{H^{3/2}(\partial\Omega)}).$
In practice, physics-informed neural networks (PINNs) are typically applied in conjunction with the choice $\mathcal{W}=H^{s}(\partial\Omega)$, $s=0$, i.e. $\mathcal{W}=L^{2}(\partial\Omega)$. However, choosing $s\neq 3/2$ means that the variational problem (13) is no longer well-posed.
In order to see this, we argue as follows. A straightforward computation reveals that
(21) $||u_{m}||_{H^{2}(\Omega)}\propto m^{3/2},\;\;\;||u_{m}||_{H^{s}(\partial\Omega)}\propto m^{s}.$
Consequently, the constant $\mathcal{C}_{2}$ in Theorem 2.1 satisfies
(22) $\mathcal{C}_{2}\leqslant||\mathcal{B}[u_{m}]||_{\mathcal{W}}^{2}/||u_{m}||_{\mathcal{X}}^{2}\propto m^{2(s-3/2)},\;\;\;\;\;\;\forall m\in\mathbb{N},$
and hence, since $m\in\mathbb{N}$ can be chosen arbitrarily large, if $s<3/2$, then $\mathcal{C}_{2}=0$. Conversely, if $s>3/2$, then a similar argument shows that the constant $\mathcal{C}_{1}$ in Theorem 2.1 satisfies
(23) $\mathcal{C}_{1}\geqslant||\mathcal{B}[u_{m}]||_{\mathcal{W}}^{2}/||u_{m}||_{\mathcal{X}}^{2}\propto m^{2(s-3/2)},\;\;\;\;\;\;\forall m\in\mathbb{N},$
in which case $\mathcal{C}_{1}$ must be unbounded.
These arguments show that the well-posedness result of Theorem 2.1 holds for the Poisson equation only if $s=3/2$, which means that the space $\mathcal{W}$ is dictated by the variational problem and cannot be chosen arbitrarily by the practitioner as is often implied in the literature.
In order to see the practical implications of using the $H^{s}(\partial\Omega)$, $s\neq 3/2$ norm, rather than the $H^{3/2}(\partial\Omega)$ norm, we return to the case where the true solution has the form $r^{m}\sin(m\theta)$ and we take advantage of (21) to replace the $H^{s}$-norm on the boundary by a weighted $L^{2}$-norm, i.e. $||u_{m}||_{H^{s}(\partial\Omega)}=m^{s}||u_{m}||_{L^{2}(\partial\Omega)}$. This is equivalent to taking $\mathcal{W}=L^{2}(\partial\Omega)$ and penalty parameter $\delta=(m^{s})^{2}=m^{2s}$. Of course, this approach will not be possible for computing the $H^{s}(\partial\Omega)$ for general problems but will suffice for the current example.
We first illustrate the scenario when $s=0$ and consider (19) with $m=2,4,8,16$. Hyperparameters for this example are provided in Appendix B. Figure 1 shows the error in various metrics for the Galerkin neural network method with formulation (19). For the $L^{2}(\Omega)$, $H^{1}(\Omega)$, and $H^{2}(\Omega)$ norms, we observe a delay in the onset of convergence as $m$ is increased. As $m$ increases, the boundary data becomes more oscillatory; the underpenalization of the boundary data for $s=0$ means that the resolution of the interior terms in (19) is prioritized in the earlier Galerkin neural network iterations at the expense of satisfying the boundary condition, resulting in the behavior observed in Figure 1.
Figure 2: Errors with respect to the number of basis functions learned using the Galerkin neural network applied to the variational problem (19) for several choices of penalty parameter $\delta=m^{2s}$ with $m=16$ and $s=0,1/2,1,3/2,2,5/2$.
In order to further amplify this point, Figure 2 shows the results when we fix $m=16$ and consider the choice $s=0,1/2,1,3/2,2,5/2$. Regardless of whether the error is measured in the $L^{2}(\Omega)$, $H^{1}(\Omega)$, or $H^{2}(\Omega)$ norms, the greatest accuracy is obtained when $s=3/2$.
The above example illustrates the importance of using the correct norms that arise in the well-posedness of the problem. In practice, the presence of fractional norms presents difficulties but is possible using techniques such as those in [21]. However, generally speaking their computation is impractical and we instead use the Sobolev norm whose order is the integer part of the order of the desired Sobolev norm. For instance, the $H^{1}(\partial\Omega)$ penalization in Figure 2 still performs quite well when compared to the optimal $H^{3/2}(\partial\Omega)$ penalization.
2.3 Extended neural network architectures
We return to the question of how to approximate the solution of variational problems of the form
(24) $u\in\mathcal{X}\;:\;a_{LS}(u,v)=F_{LS}(v)\;\;\;\forall v\in\mathcal{X},$
where $a_{LS}$ and $F_{LS}$ are defined analogously to (13). The focus of this section, in contrast to that of Section 2.2, is on problems with known solution structures.
Let $\Phi:\Omega\times\mathbb{R}\to\mathbb{R}$ be a given function and define $\mathcal{I}_{\Phi}:=\{\mu\in\mathbb{R}\;:\;\Phi(\cdot,\mu)\in\mathcal{X}\}$. For a given $m\in\mathbb{N}$, define
(25) $V_{m}^{\Phi}:=\left\{v\;:\;\Omega\to\mathbb{R}\;\bigg{|}\;v(x)=\sum_{j=1}^{m}d_{j}\Phi(x;\mu_{j}),\;\mathbf{d}\in\mathbb{R}^{m},\;\mu_{j}\in\mathcal{I}_{\Phi}\right\},$
and, for fixed $\mu\in\mathcal{I}_{\Phi}^{m}$, we also define the subset $V_{m}^{\Phi}(\mu)$ by
(26) $V_{m}^{\Phi}(\mu):=\left\{v\;:\;\Omega\to\mathbb{R}\;\bigg{|}\;v(x)=\sum_{j=1}^{m}d_{j}\Phi(x;\mu_{j}),\;\mathbf{d}\in\mathbb{R}^{m}\right\}.$
The main difference between the sets $V_{m}^{\Phi}$ and $V_{m}^{\Phi}(\mu)$ is that in the former, the nonlinear parameters $\mu$ are allowed to vary over the course of training while in the latter, they are fixed. We shall first present the extended Galerkin neural network approach on the simpler space $V_{m}^{\Phi}(\mu)$ in which the parameter $\mu$ are assumed to be known in advance before presenting the approach on the more complicated set $V_{m}^{\Phi}$ where $\mu$ is treated as a trainable parameter. As in the basic Galerkin neural network in Section 2.1, we define the set of realizations whose parameters are bounded:
@@ 2.3 Extended neural network architectures
The solutions of many classical PDEs exhibit well-known structures $\Psi(x;\lambda_{i})$, where $\lambda$ is a parameter that depends on the specific form of the domain, boundary conditions, and coefficients of the PDE. This means that for each fixed $\mathcal{M}\in\mathbb{N}$, the solution $u\in\mathcal{X}$ of such PDEs admits a representation of the form
(28) $u(x)=u_{\infty}(x)+u_{\Psi}(x;\lambda),$
where $u_{\infty}$ is the portion of the solution with high Sobolev regularity and $u_{\Psi}\in V_{\mathcal{M}}^{\Psi}(\lambda)$ the portion with low regularity. For instance, for the Poisson equation, the singular part of the solution near corners of a polygonal domain is a sum of terms of the form $\Psi(r,\theta;\lambda_{j})\propto r^{\lambda_{j}}\sin(\lambda_{j}\theta)$, where $\lambda_{j}$ depends on the interior angle at the corner and $(r,\theta)$ are polar coordinates centered at the corner.
Figure 3: Extended neural network architecture for extended Galerkin neural networks. Each basis function $\varphi^{NN}$ consists of a high-regularity component $\varphi_{\sigma}^{NN}$ which is the realization of a standard feedforward neural network and a knowledge-based function $\varphi_{\Psi}^{NN}$ which utilizes known solution structures to augment the solution.
If such knowledge-based functions are available, then it makes sense for them to be incorporated into the Galerkin neural network framework and seek
@@
(29) $\varphi_{i}^{NN}=\operatorname*{arg\,max}_{\begin{subarray}{c}v\in V_{\mathbf{n},L}^{\sigma,C_{1}}\bigoplus V_{\mathcal{M}}^{\Psi,C_{2}}(\lambda)\\ |||v|||=1\end{subarray}}\langle r(u_{i-1}),v\rangle.$
This means that the function $\varphi_{i}^{NN}$ has a structure that reflects that of the true solution in (28), namely it is the sum of a neural network function $\varphi_{\sigma,i}^{NN}\in V_{\mathbf{n},L}^{\sigma,C_{1}}$ and a function $\varphi_{\Psi,i}^{NN}\in V_{\mathcal{M}}^{\Psi,C_{2}}(\lambda)$ belonging to the “knowledge-based” space. Given an initial approximation $u_{0}=u_{\sigma,0}+u_{\Psi,0}$ to (24) with smooth part $u_{\sigma,0}$ and low-regularity part $u_{\Psi,0}\in V_{\mathcal{M}}^{\Psi}$, we define the space
(30) $S_{i}^{\Psi}:=\text{span}\{u_{\sigma,0}/|||u_{0}|||,u_{\Psi,0}/|||u_{0}|||,\varphi_{\sigma,1}^{NN},\varphi_{\Psi,1}^{NN},\dots,\varphi_{\sigma,i}^{NN},\varphi_{\Psi,i}^{NN}\},$
and the approximation $u_{i}$ to the solution of (24) is defined to be the solution of the problem
(31) $u_{i}\in S_{i}^{\Psi}\;:\;a(u,v)=F(v)\;\;\;\forall v\in S_{i}^{\Psi}.$
The solution may be written as a sum of smooth and singular parts $u_{i}=u_{\sigma,i}+u_{\Psi,i}$, where
(32) $u_{\sigma,i}\in\text{span}\{u_{\sigma,0}/|||u_{0}|||,\varphi_{\sigma,1}^{NN},\dots,\varphi_{\sigma,i}^{NN}\},\;\;\;u_{\Psi,i}\in\text{span}\{u_{\Psi,0}/|||u_{0}|||,\varphi_{\Psi,1}^{NN},\dots,\varphi_{\Psi,i}^{NN}\}.$
We briefly remark on the existence of the basis function $\varphi_{i}^{NN}$ defined as the maximizer of the weak residual in (29). The existence of $\varphi_{i}^{NN}$ is guaranteed if $V_{\mathbf{n},L}^{\sigma,C_{1}}\oplus V_{\mathcal{M}}^{\Psi,C_{2}}(\lambda)$ is a compact subset of $\mathcal{X}$. The compactness of $V_{\mathbf{n},L}^{\sigma,C_{1}}$ is addressed in [19, Proposition 3.5] and discussed in the context of the basic Galerkin neural network approach in [1, Remark 2.4]. To establish compactness of $V_{\mathcal{M}}^{\Psi}(\lambda)$, we utilize the following lemma.